# The radius of statistical efficiency

### v1

@@ 2 The distance to ill-conditioned problems
Moreover, if for some $c,q>0$, the inequality ${\rm RSE}(\nu)^{q}\leq c\cdot{\rm REG}(\nu)^{-1}$ holds for all $\nu\in\mathcal{Q}^{\prime}\setminus\mathcal{E}$, then the supremum in (12) is upper bounded by $c\cdot({\rm RSE}(\mu)-r)^{-1/q}$.
@@
At first sight, it appears that computing ${\rm RSE}(\mathsf{D})$ in concrete problems is difficult. Indeed, the set of ill-conditioned distributions $\mathcal{E}$ may be quite exotic and computing ${\rm RSE}(\mathsf{D})$ amounts to estimating the Wasserstein-2 distance $W_{2}(\mathsf{D},\mathcal{E})$. In contrast, computing the regularity modulus ${\rm REG}(\mathsf{D})$ should be relatively straightforward. The key observation now is that the two quantities, ${\rm RSE}(\mathsf{D})$ and ${\rm REG}(\mathsf{D})$, are closely related since $\mathcal{E}$ is the set of minimizers of the function $\mathcal{J}(\mu):=1/{\rm REG}(\mu)$. Thus it would be ideal if there were a quantitative relationship of the form:
$(W_{2}(\mu,\mathcal{E}))^{\ell_{1}}\lesssim\mathcal{J}(\mu)\lesssim(W_{2}(\mu,\mathcal{E}))^{\ell_{2}}\qquad\forall\mu\in\mathcal{P}_{2}.$ (13)
@@
The upper should be elementary to establish because it amounts to upper bounding the growth of the functional $\mathcal{J}(\mu)$. The lower bound is more substantial because it requires lower-bounding the growth of $\mathcal{J}(\mu)$ by a nonnegative function of the distance. Such lower-estimates are called error bounds in nonlinear analysis and can be checked by various “slope”-based conditions. See for example the authoritative monographs on the subject, [20, Chapter 8] and [38, Chapter 3]. Indeed, Demmel’s original work [28] relies on verifying an error bound property as well, albeit in the much simpler Euclidean setting. The following theorem provides a sufficient condition (14) ensuring the relationship (13). We state the theorem loosely by compressing all multiplicative numerical constants via the symbol $\lesssim$. More precise and sharper guarantees appear in Appendix B.2.111In the theorem statement, the symbol $DF(x)\colon\mathbf{R}^{d}\to\bm{S}^{k}$ denotes the differential of $F$, while the symbol $DF(x)^{*}\colon\bm{S}^{k}\to\mathbf{R}^{d}$ is the adjoint linear map of $DF(x)$.
@@
Theorem 2.2: (Infinitesimal characterization of RSE) Consider the problem $\mathtt{SP}(\mathsf{D})$ of supervised learning and let ${\beta}_{\star}$ be its minimizer. Set $\mathcal{Q}=\mathcal{P}_{2}$ and suppose that the Hessian $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ corresponding to a measure $\mu\in\mathcal{P}_{2}$ can be written as $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})=\operatorname{\mathbb{E}}_{\mu}F(x),$ for some $C^{1}$-smooth map $F\colon\mathbf{R}^{d}\to\bm{S}^{k}_{+}$ satisfying $\|DF(x)\|_{\rm op}\lesssim 1+\|x\|_{2}$ for all $x\in\mathbf{R}^{d}$. Suppose that there exists $q_{1},q_{2}\in[0,1)$ such that for all measures $\nu\in\mathcal{P}_{2}\setminus\mathcal{E}$, the estimate $\lambda_{\min}\left(\operatorname{\mathbb{E}}_{\nu}F(x)\right)^{q_{1}}\lesssim\sqrt{\operatorname{\mathbb{E}}_{\nu}\|DF(x)^{*}[uu^{\top}]\|^{2}_{2}}\lesssim\lambda_{\min}\left(\operatorname{\mathbb{E}}_{\nu}F(x)\right)^{q_{2}},$ (14) holds for some eigenvector $u\in\mathbb{S}^{d-1}$ of the matrix $\operatorname{\mathbb{E}}_{\nu}F(x)$ corresponding to its minimal eigenvalue. Then for every $\mu\in\mathcal{P}_{2}$ the inequality holds: ${\rm REG}(\mu)^{q_{2}-1}\lesssim{\rm RSE}(\mu)\lesssim{\rm REG}(\mu)^{q_{1}-1}.$ (15)
@@
$\left(\operatorname{\mathbb{E}}_{\nu}\langle u,g(x)\rangle^{2}\right)^{q_{1}}\lesssim\sqrt{\operatorname{\mathbb{E}}_{\nu}\langle u,g(x)\rangle^{2}\|\nabla g(x)u]\|^{2}_{2}}\lesssim\left(\operatorname{\mathbb{E}}_{\nu}\langle u,g(x)\rangle^{2}\right)^{q_{2}},$
Observe that all three terms would match exactly with $q_{1}=q_{2}=\frac{1}{2}$ were it not for the term $\|\nabla g(x)u]\|^{2}_{2}$ that reweighs the middle integral. It is this reweighing that may impact the values of $q_{1}$ and $q_{2}$. The salient feature of Theorem 2 is that it completely circumvents the need for explicitly estimating the distance to the exceptional set $\mathcal{E}$. One could apply this theorem to a number of examples that will appear in the rest of the paper. That being said, in all the upcoming examples, we will be able to compute the distance to $\mathcal{E}$ explicitly, thereby obtaining sharper estimates than would follow from Theorem 2. Nonetheless, we believe that Theorem 2 is interesting in its own right and may be useful for analyzing RSE in more complex situations.
@@ 3 Principal component analysis
Principal component analysis (PCA) is a common technique for dimension reduction. The goal of PCA is to find a low-dimensional subspace that captures the majority of the variance of the distribution. In this section, we compute the radius of statistical efficiency for PCA. Setting the stage, let $x$ be a random vector in $\mathbf{R}^{d}$ drawn from a zero-mean distribution $\mathsf{D}\in\mathcal{P}_{2}^{\circ}(\mathbf{R}^{d})$. A unit vector $v$ for which the random variable $\langle v,x\rangle$ has maximal variance is called the first principal component of $\mathsf{D}$. Thus, the first principal component is the maximizer of the problem
$\max_{v\in\mathbb{S}^{d-1}}~{}\tfrac{1}{2}\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathsf{D}}~{}\langle v,x\rangle^{2},$ (16)
Equivalently, the first principal component is the eigenvector corresponding to the maximal eigenvalue of the covariance matrix $\Sigma_{\mathsf{D}}=\operatorname{\mathbb{E}}_{\mathsf{D}}xx^{\top}$. Intuitively, the problem (16) is more challenging when the gap between the top two eigenvalues of $\Sigma_{\mathsf{D}}$ is small. This is the content of the following lemma, whose proof appears in Appendix C.1.
@@
=$\mathcal{E}=\{\mu\in\mathcal{P}_{2}^{\circ}:\lambda_{1}(\Sigma_{\mu})=\lambda_{2}(\Sigma_{\mu})\}.$
Moreover, for any $\mu\in\mathcal{P}_{2}^{\circ}\setminus\mathcal{E}$, equality ${\rm REG}(\mu)^{-1}=\lambda_{1}(\Sigma_{\mu})-\lambda_{2}(\Sigma_{\mu})$ holds.
Therefore, estimating RSE amounts to computing the Wasserstein-2 distance of a base measure $\mathsf{D}$ to the set $\mathcal{E}$. The end result is the following theorem; we defer its proof to Appendix C.2.
Theorem 3.2: (RSE for top principal component) Consider the problem (16) and define the covariance $\Sigma_{\mathsf{D}}:=\operatorname{\mathbb{E}}_{\mathsf{D}}[xx^{\top}]$. Then, equality holds: $\operatorname{RSE}(\mathsf{D})=\tfrac{1}{\sqrt{2}}\left(\sqrt{\lambda_{1}(\Sigma_{\mathsf{D}})}-\sqrt{\lambda_{2}(\Sigma_{\mathsf{D}})}\right).$ (17) In particular, we have $\operatorname{RSE}(\mathsf{D})\cdot{\rm REG}(\mathsf{D})=\frac{1}{\sqrt{2}\left(\sqrt{\lambda_{1}(\Sigma_{\mathsf{D}})}+\sqrt{\lambda_{2}(\Sigma_{\mathsf{D}})}\right)}.$ (18)
Thus treating $\lambda_{1}(\Sigma_{\mathsf{D}})$ in (18) as being of constant order, we see that the hardness of the problem is inversely proportional to the distance to the nearest ill-posed problem, ${\rm REG}(\mathsf{D})\propto\operatorname{RSE}(\mathsf{D})^{-1}$.
More generally still, we may be interested in finding a $q$-dimensional subspace $\mathcal{V}\subset\mathbf{R}^{d}$ that captures most of the variance. Analytically, this amounts to solving the problem
$\max_{R\in{\rm Gr}(q,d)}~{}f(R)=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathsf{D}}\|Rx\|^{2}_{2},$ (19)
where the Grassmannian manifold ${\rm Gr}(q,d)$ consists of all orthogonal projections $R\in\bm{S}^{d}$ onto $q$ dimensional subspaces of $\mathbf{R}^{d}$. The column space of the optimal matrix $R$ is called the top $q$ principal subspace. Equivalently, the top $q$ principal subspace is the span of the eigenspaces of $\Sigma_{\mathsf{D}}$ corresponding to its top $q$ eigenvalues. The following lemma is a direction extension of Lemma 3.1; see Appendix C.3 for a proof.
@@
=$\mathcal{E}=\{\mu\in\mathcal{P}_{2}^{\circ}:\lambda_{q}(\Sigma_{\mu})=\lambda_{q+1}(\Sigma_{\mu})\}.$
Moreover, for any $\mu\in\mathcal{P}_{2}^{\circ}\setminus\mathcal{E}$, equality ${\rm REG}(\mu)^{-1}=\lambda_{q}(\Sigma_{\mu})-\lambda_{q+1}(\Sigma_{\mu})$ holds.
Thus, estimating RSE amounts to computing the Wasserstein-2 distance of a base measure $\mathsf{D}$ to the set $\mathcal{E}$. The end result is the following theorem; the proof appears in Appendix C.4.
Theorem 3.4: (RSE for PCA) Consider the problem (19) and define the covariance $\Sigma:=\operatorname{\mathbb{E}}_{\mathsf{D}}[xx^{\top}]$. Then, the estimate holds: $\operatorname{RSE}(\mathsf{D})=\tfrac{1}{\sqrt{2}}\left(\sqrt{\lambda_{q}(\Sigma)}-\sqrt{\lambda_{q+1}(\Sigma)}\right).$ In particular, we have $\operatorname{RSE}(\mathsf{D})\cdot{\rm REG}(\mathsf{D})=\frac{1}{\sqrt{2}\left(\sqrt{\lambda_{q}(\Sigma_{\mathsf{D}})}+\sqrt{\lambda_{q+1}(\Sigma_{\mathsf{D}})}\right)}.$ (20)
Thus similarly to the case $q=1$, treating $\lambda_{q}(\Sigma_{\mathsf{D}})$ in (20) as being of constant order, the hardness of the problem is inversely proportional to the distance to ill-posed problems, ${\rm REG}(\mathsf{D})\propto\operatorname{RSE}(\mathsf{D})^{-1}$.
The proofs of Theorems 3 and 3 rely on estimating the distance to the exceptional sets $\mathcal{E}$. Notice that these two are defined purely in terms of the spectrum of the second-moment matrix. Although such “spectral” sets in $\mathcal{P}_{2}$ are quite complicated, their distance can be readily computed. Geometric properties of such sets are explored in Appendix A.2 and may be of independent interest.
@@ 4 Generalized linear models and (quasi) maximum likelihood estimation
In this section, we compute the RSE for a large class of supervised learning problems arising from (quasi) maximum likelihood estimation (QMLE). The goal is to estimate a parameter ${\beta}_{\star}\in\mathcal{M}$ where the constraint set $\mathcal{M}$ is a $C^{2}$-smooth embedded submanifold of $\mathbf{R}^{d}$. Setting the stage, suppose that we have an $L^{2}$ random vector $x$ (the predictor) and an $L^{2}$ random variable $y$ (the response) satisfying the GLM conditions
${\operatorname{\mathbb{E}}}_{\mathsf{D}}\big{[}y\!\mid\!x\big{]}=h^{\prime}(\langle x,\beta_{\star}\rangle)\qquad\text{and}\qquad{\operatorname{Var}}_{\mathsf{D}}\big{[}y\!\mid\!x\big{]}=\sigma^{2}\cdot h^{\prime\prime}(\langle x,\beta_{\star}\rangle)$ (21)
[…]
@@
$\min_{\beta\in\mathcal{M}}f(\beta):={\operatorname{\mathbb{E}}}_{\mathsf{D}}\big{[}h(\langle x,\beta\rangle)-y\langle x,\beta\rangle\big{]}.$ (22)
The function $f$ given in (22) is called the negative log quasi-likelihood of the GLM (21). In general $f$ is not the negative of the log likelihood function, yet it shares many of its properties and hence the name. The motivation for this loss function comes from the canonical example of (21), where the conditional density of $y$ given $x$ admits an exponential-family formulation. In this case, standard maximum likelihood estimation of ${\beta}_{\star}$ coincides with (22). As illustration, Table 1 lists some common examples of QMLE. Henceforth, we let $\mathcal{T}:=T_{\mathcal{M}}({\beta}_{\star})$ denote the tangent space of $\mathcal{M}$ at ${\beta}_{\star}$.
@@
Linear $y=\langle x,{\beta}_{\star}\rangle+\varepsilon$ $\tfrac{1}{2}\theta^{2}$ $1$
Logistic $y\!\mid\!x\sim\mathsf{Ber}\Big{(}\tfrac{\exp\langle x,{\beta}_{\star}\rangle}{1+\exp\langle x,{\beta}_{\star}\rangle}\Big{)}$ $\log(1+\exp\theta)$ $\frac{\exp(\theta)}{(1+\exp(\theta))^{2}}$
Poisson $y\!\mid\!x\sim\mathsf{Poi}({\exp}\langle x,{\beta}_{\star}\rangle)$ $\exp\theta$ $\exp(\theta)$
Gamma $y\!\mid\!x\sim\Gamma(\sigma^{-2},-\sigma^{-2}\langle x,{\beta}_{\star}\rangle)$ $-\log(-\theta)$ $\theta^{-2}$
Table 1: Examples of QGLM problems (21). In the linear regression model, $\varepsilon$ is random noise satisfying ${\operatorname{\mathbb{E}}}[\varepsilon\!\mid\!x]=0$ and ${\operatorname{Var}}[\varepsilon\!\mid\!x]=\sigma^{2}$.
We begin with the following lemma that characterizes the set of ill-conditioned problem instances.
@@
=$\mathcal{E}=\left\{\mu\in\mathcal{Q}:\mathcal{T}\cap\ker\left(\Sigma_{\mu}\right)\neq\{0\}\right\}.$
@@
Moreover, for any $\mu\in\mathcal{Q}\setminus\mathcal{E}$ for which there exist $c_{\texttt{lb}},c_{\texttt{ub}}>0$ satisfying $c_{\texttt{lb}}\leq h^{\prime\prime}(\langle x,\beta_{\star}\rangle)\leq c_{\texttt{ub}}$ for $\mu$-almost every $x$, we have $c_{\texttt{lb}}\leq{\rm REG}(\mu)\cdot\lambda_{\min}\left(\Sigma_{\mu}|_{{\mathcal{T}}}\right)\leq c_{\texttt{ub}}$.
@@
The proof of this lemma is deferred to Appendix D.1. The RSE for the problem follows quickly by computing the distance to the set $\mathcal{E}$ in Lemma 4.1; see Appendix D.2 for a proof.

### v2

@@ 2 The distance to ill-conditioned problems
Moreover, if for some $c,q>0$, the inequality ${\rm RSE}(\nu)^{q}\leq c\cdot{\rm REG}(\nu)^{-1}$ holds for all $\nu\in\mathcal{Q}^{\prime}\setminus\mathcal{E}$, then the supremum in (12) is upper-bounded by $c\cdot({\rm RSE}(\mu)-r)^{-q}$.
@@
At first sight, it appears that computing ${\rm RSE}(\mathsf{D})$ in concrete problems is difficult. Indeed, the set of ill-conditioned distributions $\mathcal{E}$ may be quite exotic and computing ${\rm RSE}(\mathsf{D})$ amounts to estimating the Wasserstein-2 distance $W_{2}(\mathsf{D},\mathcal{E})$. In contrast, computing the regularity modulus ${\rm REG}(\mathsf{D})$ should be relatively straightforward. The key observation now is that the two quantities ${\rm RSE}(\mathsf{D})$ and ${\rm REG}(\mathsf{D})$ are closely related since $\mathcal{E}$ is the set of minimizers of the function $\mathcal{J}(\mu):={\rm REG}(\mu)^{-1}$ (i.e., $\mathcal{E}$ coincides with the zero set $\mathcal{J}=0]$). Thus, it would be ideal if there were a quantitative relationship of the form
$(W_{2}(\mu,\mathcal{E}))^{\ell_{1}}\lesssim\mathcal{J}(\mu)\lesssim(W_{2}(\mu,\mathcal{E}))^{\ell_{2}}\qquad\forall\mu\in\mathcal{P}_{2}$ (13)
@@
for some constants $\ell_{1},\ell_{2}\geq 1$. The upper bound should be elementary to establish because it amounts to upper-bounding the growth of the function $\mathcal{J}(\mu)$. The lower bound is more substantial because it requires lower-bounding the growth of $\mathcal{J}(\mu)$ by a nonnegative function of the distance. Such lower-estimates are called error bounds in nonlinear analysis and can be checked by various “slope”-based conditions. See, for example, the authoritative monographs on the subject [20, Chapter 8] and [39, Chapter 3]. Indeed, Demmel’s original work [29] relies on verifying an error bound property as well, albeit in the much simpler Euclidean setting. The following theorem provides a sufficient condition (14) ensuring the relationship (13), which appears in the equivalent form
${\rm REG}(\mu)^{q_{2}-1}\lesssim{\rm RSE}(\mu)\lesssim{\rm REG}(\mu)^{q_{1}-1},$
where $q_{i}\in[0,1)$ corresponds to (13) via $q_{i}=1-\ell_{i}^{-1}$ for $i=1,2$. We state the theorem loosely by compressing all multiplicative numerical constants via the symbol $\lesssim$. More precise and sharper guarantees appear in Appendix B.2. In the theorem statement, the symbol $DF(x)\colon\mathbf{R}^{d}\to\bm{S}^{k}$ denotes the differential of $F$ at $x$, while the symbol $DF(x)^{*}\colon\bm{S}^{k}\to\mathbf{R}^{d}$ denotes the adjoint linear map of $DF(x)$.
@@
Theorem 2.2: (Infinitesimal characterization of RSE) Consider the problem $\mathtt{SP}(\mathsf{D})$ of supervised learning and let $\beta$$_{\star}$ be its minimizer. Set $\mathcal{Q}=\mathcal{P}_{2}$ and suppose the Hessian $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ corresponding to any measure $\mu\in\mathcal{P}_{2}$ can be written as $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})={\operatorname{\mathbb{E}}}_{\mu}F(x)$ for some $C^{1}$-smooth map $F\colon\mathbf{R}^{d}\to\bm{S}^{k}_{+}$ satisfying $\|DF(x)\|_{\rm op}\lesssim 1+\|x\|_{2}$ for all $x\in\mathbf{R}^{d}$. Suppose there exist $q_{1},q_{2}\in[0,1)$ such that for all measures $\nu\in\mathcal{P}_{2}\setminus\mathcal{E}$, the estimate $\lambda_{\min}({\operatorname{\mathbb{E}}}_{\nu}F(x))^{q_{1}}\lesssim\sqrt{{\operatorname{\mathbb{E}}}_{\nu}\|DF(x)^{*}[uu^{\top}]\|^{2}_{2}}\lesssim\lambda_{\min}({\operatorname{\mathbb{E}}}_{\nu}F(x))^{q_{2}}$ (14) holds for some eigenvector $u\in\mathbb{S}^{k-1}$ of the matrix ${\operatorname{\mathbb{E}}}_{\nu}F(x)$ corresponding to its minimal eigenvalue. Then for every $\mu\in\mathcal{P}_{2}$, the following estimate holds: ${\rm REG}(\mu)^{q_{2}-1}\lesssim{\rm RSE}(\mu)\lesssim{\rm REG}(\mu)^{q_{1}-1}.$ (15)
@@
$\left(\operatorname{\mathbb{E}}_{\nu}\langle u,g(x)\rangle^{2}\right)^{q_{1}}\lesssim\sqrt{{\operatorname{\mathbb{E}}}_{\nu}\Big[\langle u,g(x)\rangle^{2}\|\nabla g(x)^{\top}u\|^{2}_{2}\Big]}\lesssim\left(\operatorname{\mathbb{E}}_{\nu}\langle u,g(x)\rangle^{2}\right)^{q_{2}}.$
Observe that all three terms would match exactly with $q_{1}=q_{2}=\frac{1}{2}$ were it not for the term $\|\nabla g(x)^{\top}u\|^{2}_{2}$ that reweighs the middle integral. It is this reweighing that may impact the values of $q_{1}$ and $q_{2}$. The salient feature of Theorem 2 is that it completely circumvents the need for explicitly estimating the distance to the exceptional set $\mathcal{E}$. One could apply this theorem to some of the examples that will appear in the rest of the paper. That being said, in all the upcoming examples, we will be able to compute the distance to $\mathcal{E}$ explicitly, thereby obtaining sharper estimates than would follow from Theorem 2. The main drawback of this result is its stringent assumptions on the differential of $F,$ which may limit its applicability. Nonetheless, we believe that Theorem 2 is interesting in its own right and may be useful for analyzing RSE in certain situations.
@@ 3 Principal component analysis
Principal component analysis (PCA) is a common technique for dimension reduction. The goal of PCA is to find a low-dimensional subspace that captures the majority of the variance of the distribution. In this section, we compute the radius of statistical efficiency for PCA. Setting the stage, let $x$ be a random vector in $\mathbf{R}^{d}$ drawn from a zero-mean distribution $\mathsf{D}\in\mathcal{P}_{2}^{\circ}(\mathbf{R}^{d})$. A unit vector $v$ for which the random variable $\langle v,x\rangle$ has maximal variance is called the first principal component of $\mathsf{D}$. Thus, the first principal component is a maximizer of the problem
$\max_{v\in\mathbb{S}^{d-1}}\,\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathsf{D}}\tfrac{1}{2}\langle v,x\rangle^{2}.$ (16)
Equivalently, the first principal component is a unit eigenvector corresponding to the maximal eigenvalue of the covariance matrix $\Sigma_{\mathsf{D}}:=\operatorname{\mathbb{E}}_{\mathsf{D}}[xx^{\top}]$. Intuitively, the problem (16) is more challenging when the gap between the top two eigenvalues of $\Sigma_{\mathsf{D}}$ is small. This is the content of the following lemma, whose proof appears in Appendix C.1.
@@
$\mathcal{E}=\{\mu\in\mathcal{P}_{2}^{\circ}\mid\lambda_{1}(\Sigma_{\mu})=\lambda_{2}(\Sigma_{\mu})\}.$
Moreover, for any $\mu\in\mathcal{P}_{2}^{\circ}$, equality ${\rm REG}(\mu)^{-1}=\lambda_{1}(\Sigma_{\mu})-\lambda_{2}(\Sigma_{\mu})$ holds.
Computing the RSE amounts to determining the Wasserstein-2 distance of the base measure $\mathsf{D}$ to the set $\mathcal{E}$. The end result is the following theorem; we defer its proof to Appendix C.2.
Theorem 3.2: (RSE for top principal component) Consider the problem (16) with covariance matrix $\Sigma_{\mathsf{D}}:=\operatorname{\mathbb{E}}_{\mathsf{D}}[xx^{\top}]$. Then equality holds: $\operatorname{RSE}(\mathsf{D})=\tfrac{1}{\sqrt{2}}\bigg(\sqrt{\lambda_{1}(\Sigma_{\mathsf{D}})}-\sqrt{\lambda_{2}(\Sigma_{\mathsf{D}})}\bigg).$ (17) In particular, if $\mathsf{D}\in\mathcal{P}_{2}^{\circ}\setminus\mathcal{E}$, then $\operatorname{RSE}(\mathsf{D})\cdot{\rm REG}(\mathsf{D})=\frac{1}{\sqrt{2}\Big(\sqrt{\lambda_{1}(\Sigma_{\mathsf{D}})}+\sqrt{\lambda_{2}(\Sigma_{\mathsf{D}})}\Big)}.$ (18)
More generally, we may be interested in finding a $q$-dimensional subspace $\mathcal{V}\subset\mathbf{R}^{d}$ that captures most of the variance. Analytically, this amounts to solving the problem
$\max_{R\in{\rm Gr}(q,d)}~f(R):=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathsf{D}}\|Rx\|^{2}_{2},$ (19)
where the Grassmannian manifold ${\rm Gr}(q,d)$ consists of all orthogonal projections $R\in\bm{S}^{d}$ onto $q$-dimensional subspaces of $\mathbf{R}^{d}$. The column space of an optimal matrix $R$ is called the top $q$ principal subspace. Equivalently, the top $q$ principal subspace is the span of the eigenvectors of $\Sigma_{\mathsf{D}}$ corresponding to its top $q$ eigenvalues. The following lemma is a direct extension of Lemma 3.1; see Appendix C.3 for a proof.
@@
$\mathcal{E}=\{\mu\in\mathcal{P}_{2}^{\circ}\mid\lambda_{q}(\Sigma_{\mu})=\lambda_{q+1}(\Sigma_{\mu})\}.$
Moreover, for any $\mu\in\mathcal{P}_{2}^{\circ}$, equality ${\rm REG}(\mu)^{-1}=\lambda_{q}(\Sigma_{\mu})-\lambda_{q+1}(\Sigma_{\mu})$ holds.
Again, computing the RSE amounts to determining the Wasserstein-2 distance of the base measure $\mathsf{D}$ to the set $\mathcal{E}$. The end result is the following theorem; the proof appears in Appendix C.4.
Theorem 3.4: (RSE for PCA) Consider the problem (19) with covariance matrix $\Sigma_{\mathsf{D}}:=\operatorname{\mathbb{E}}_{\mathsf{D}}[xx^{\top}]$. Then equality holds: $\operatorname{RSE}(\mathsf{D})=\tfrac{1}{\sqrt{2}}\bigg(\sqrt{\lambda_{q}(\Sigma_{\mathsf{D}})}-\sqrt{\lambda_{q+1}(\Sigma_{\mathsf{D}})}\bigg).$ In particular, if $\mathsf{D}\in\mathcal{P}_{2}^{\circ}\setminus\mathcal{E}$, then $\operatorname{RSE}(\mathsf{D})\cdot{\rm REG}(\mathsf{D})=\frac{1}{\sqrt{2}\Big(\sqrt{\lambda_{q}(\Sigma_{\mathsf{D}})}+\sqrt{\lambda_{q+1}(\Sigma_{\mathsf{D}})}\Big)}.$ (20)
We see two different regimes for the reciprocal relationship between RSE and REG. When
$\sqrt{\lambda_{q}(\Sigma_{\mathsf{D}})}+\sqrt{\lambda_{q+1}(\Sigma_{\mathsf{D}})}\gg\sqrt{\lambda_{q}(\Sigma_{\mathsf{D}})}-\sqrt{\lambda_{q+1}(\Sigma_{\mathsf{D}})},$ (21)
[…]
The proofs of Theorems 3 and 3 rely on estimating the distance to the exceptional sets $\mathcal{E}$. Notice that these two sets are defined purely in terms of the spectrum of the covariance matrix. Although such “spectral” sets in $\mathcal{P}_{2}^{\circ}$ are quite complicated, their distance can be readily computed. Geometric properties of such sets are explored in Appendix A.2 and may be of independent interest.
@@ 4 Generalized linear models and (quasi) maximum likelihood estimation
In this section, we compute the RSE for a large class of supervised learning problems arising from (quasi) maximum likelihood estimation (QMLE). The goal is to estimate a parameter $\beta$$_{\star}\in\mathcal{M}$ where the constraint set $\mathcal{M}$ is a $C^{2}$-smooth embedded submanifold of $\mathbf{R}^{d}$. Setting the stage, suppose that we have an $L^{2}$ random feature vector $x$ and an $L^{2}$ random label variable $y$ satisfying the generalized linear model (GLM) conditions
${\operatorname{\mathbb{E}}}\big[y\!\mid\!x\big]=h^{\prime}(\langle x,\beta_{\star}\rangle)\qquad\text{and}\qquad{\operatorname{Var}}\big[y\!\mid\!x\big]=\sigma^{2}\cdot h^{\prime\prime}(\langle x,\beta_{\star}\rangle)$ (22)
for some known $C^{2}$-smooth convex function $h\colon\mathbf{R}\rightarrow\mathbf{R}$ with $h^{\prime\prime}>0$ and parameter $\sigma^{2}>0$. The function $h$ is called the cumulant function of the model (22), and $\sigma^{2}$ the dispersion parameter. Here and from now on, we make the blanket assumption that the distribution $\mathsf{D}_{x}$ of $x$ lies in a subset $\mathcal{Q}$ of $\mathcal{P}_{2}$ comprised of probability measures satisfying sufficient regularity conditions to take expectations and to interchange differentiation and expectation as necessary. In particular, we assume that the function $\phi\colon\mathcal{M}\rightarrow\mathbf{R}$ given by $\phi(\beta)={\operatorname{\mathbb{E}}}[h(\langle x,\beta\rangle)]$ is well defined and has a $C^{2}$-smooth local extension to a neighborhood of $\beta$$_{\star}$ in $\mathbf{R}^{d}$ with $\nabla\phi({\beta}_{\star})={\operatorname{\mathbb{E}}}[h^{\prime}(\langle x,{\beta}_{\star}\rangle)x]$ and $\nabla^{2}\phi({\beta}_{\star})=\operatorname{\mathbb{E}}[h^{\prime\prime}(\langle x,{\beta}_{\star}\rangle)xx^{\top}]$.
@@
$\min_{\beta\in\mathcal{M}}f(\beta):=\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\mathsf{D}}\big[h(\langle x,\beta\rangle)-y\langle x,\beta\rangle\big].$ (23)
The function $f$ given in (23) is called the negative log quasi-likelihood of the GLM (22). In general, $f$ is not the negative of the log likelihood function, yet it shares many of its properties and hence the name. The motivation for this loss function comes from the canonical example of (22) where the conditional density of $y$ given $x$ admits an exponential-family formulation. In this case, standard maximum likelihood estimation of $\beta$$_{\star}$ coincides with (23). As an illustration, Table 1 lists some common GLM examples.
@@
Linear $y=\langle x,{\beta}_{\star}\rangle+\varepsilon$ $\tfrac{1}{2}\theta^{2}$ 11
Logistic $y\!\mid\!x\sim\mathsf{Ber}\Big(\tfrac{\exp\langle x,{\beta}_{\star}\rangle}{1+\exp\langle x,{\beta}_{\star}\rangle}\Big)$ $\log(1+\exp\theta)$ $\frac{\exp\theta}{(1+\exp\theta)^{2}}$
Poisson $y\!\mid\!x\sim\mathsf{Poi}({\exp}\langle x,{\beta}_{\star}\rangle)$ $\exp\theta$ $\exp\theta$
Gamma $y\!\mid\!x\sim\Gamma(\sigma^{-2},-\sigma^{-2}\langle x,{\beta}_{\star}\rangle)$ $\log(-\theta)$ $\theta^{-2}$
Table 1: GLM examples corresponding to (22). In the linear regression model, $\varepsilon$ is random noise satisfying ${\operatorname{\mathbb{E}}}[\varepsilon\!\mid\!x]=0$ and ${\operatorname{Var}}[\varepsilon\!\mid\!x]=\sigma^{2}$.
We begin with the following lemma that characterizes the set of ill-conditioned problem instances. Henceforth, we let $\mathcal{T}:=T_{\mathcal{M}}({\beta}_{\star})$ denote the tangent space of $\mathcal{M}$ at $\beta$$_{\star}$.
@@
$\mathcal{E}=\big\{\mu\in\mathcal{Q}\mid\mathcal{T}\cap\ker(\Sigma_{\mu})\neq\{0\}\big\}.$
@@
Moreover, for any $\mu\in\mathcal{Q}\setminus\mathcal{E}$ for which there exist $c_{\textup{{lb}}},c_{\textup{{ub}}}>0$ satisfying $c_{\textup{{lb}}}\leq{h^{\prime\prime}(\langle x,\beta_{\star}\rangle)^{-1}}\leq c_{\textup{{ub}}}$ for $\mu$-almost every $x$, we have $c_{\textup{{lb}}}\leq{\rm REG}(\mu)\cdot\lambda_{\min}(\Sigma_{\mu}|_{{\mathcal{T}}})\leq c_{\textup{{ub}}}$.
@@
The proof of this lemma is deferred to Appendix D.1. The RSE for the problem (23) follows quickly by computing the distance to the set $\mathcal{E}$ in Lemma 4.1; see Appendix D.2 for a proof.